OpenStudy (anonymous):

y=-2-2x, y=-1/2+1 whATS THE SOLUTION FOR THIS GRAPH

OpenStudy (mathmagician):

from the second equation, you get y=-0.5+1=0.5 Then the first equation becomes 0.5=-2-2x and therefore x=-1.25 so the answer is (x;y)=(-1.25;0.5)

OpenStudy (anonymous):

(-5/4,1/2) thanks :-)